OpenStudy (anonymous):

Use a trigonometric ratio to find the value of x. Round your answer to the nearest tenth. The triangle is not drawn to scale. Need Help! Please I don't get it at all I tried to solve it.

OpenStudy (anonymous):

Post the graph and I will try

OpenStudy (anonymous):

Just give me one sec

OpenStudy (anonymous):

Here it is

OpenStudy (anonymous):

Are you still there Jacich?

OpenStudy (anonymous):

HI Welcome to Openstudy! Do you know the trigonometric ratios? SOH CAH TOA? \(\sf \Large sin\ \theta= \frac{opposite}{adjacent}\\ \Large cos\ \theta= \frac{adjacent}{hypotenuse} \\ \Large tan\ \theta = \frac{opposite}{adjacent} \) Now look at the image below:

OpenStudy (anonymous):

You are given the \(\sf \theta = 20°\ and\ opposite=9\) You are looking for the adjacent side, which is x. so which of the trig ratios can you use?

OpenStudy (anonymous):

**just refresh your browser of you see question marks on you screen

OpenStudy (anonymous):

What do you mean by trig ratios? Sorry this is new to me. Can explain is step by step. Please if it's not to much trouble.

OpenStudy (anonymous):

no worries! we're here to help I mentioned them above :)

OpenStudy (anonymous):

Sin and tan?

OpenStudy (anonymous):

Trig ratios are SOH CAH TOA (just abbreviated form) their ratios are mentioned above and I labelled your diagram: 

OpenStudy (anonymous):

since you are given the opposite and the theta, and you are looking for x or the adjacent side, you can use \(\sf \Large tan \theta = \frac{opposite}{adjacent}\) just plug in your given values

OpenStudy (anonymous):

so would it be like 20/9?

OpenStudy (anonymous):

\(\sf \Large tan\ 20= \frac{9}{adjacent}\) evaluate this, it will be: \(\sf \Large \frac{tan\ 20}{9}= adjacent\)

OpenStudy (anonymous):

so would you just divide\[\tan 20 \div 9=adj\] which would give you .45?

OpenStudy (anonymous):

oops my bad! I'm sorry! it should be \(\sf \Large \frac{9}{tan 20}= adjacent\)

OpenStudy (anonymous):

So it would be .45 right?

OpenStudy (anonymous):

nope. you should get a different number because that one is too small.

OpenStudy (anonymous):

recalculate it again and let's see if we get the same

OpenStudy (anonymous):

2.22?

OpenStudy (anonymous):

no I got a different one be careful when plugging the numbers in your calculator

OpenStudy (anonymous):

ok, for what the calculator is telling me 9/20= 0.45 or 20/9= 2.22

OpenStudy (anonymous):

you forgot to enter the "tan" function

OpenStudy (anonymous):

it should look like this:

OpenStudy (anonymous):

Oops my bad. should it equal to 4.02?

OpenStudy (anonymous):

yes kind of. your calculator is set in radian mode, it's better if you change it to degrees because your theta is in degrees. go to the mode of your calculator and click on the 'deg' mode. then calculate it again.

OpenStudy (anonymous):

Just give me one sec I will be right back. But it is still giving me 0.45

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

Ok Im back and I did it again and it gave me 4.02295598054

OpenStudy (anonymous):

oh wait!! I think I'm the one who made a mistake here You are right (: I also got 4.02 lol i'm getting 24 in a different calculator, but then I use my own and I got the same thing that you get :3 again i'm sorry, I hope this will not stop you to stay here at openstudy

OpenStudy (anonymous):

Are you still there Data_LG2?

OpenStudy (anonymous):

Its ok, so what would be the ratio to the problem?